(1/2)^4=1/16 write in logarithmic form
Answer by stanbon(57967)

(1/2)^4=1/16 write in logarithmic form
The base is (1/2)
The exponent is 4
The result is 1/16
Since log means exponent you write:
log(base 1/2) 1/16 = 4
cheers,
Stan H.